OpenStudy (princess123):

what is 5√72 in simplest radical form?

OpenStudy (anonymous):

30\[30\sqrt{2}\]

OpenStudy (anonymous):

\[5*\sqrt{36*2}=5*6\sqrt{2}=30\sqrt{2}\]

OpenStudy (princess123):

thanks!

OpenStudy (anonymous):

welcm :)